OpenStudy (anonymous):

Find the variables 4w - 2x = 4 + x 12 + 3x = 7 - 6w

OpenStudy (imstuck):

I'm assuming the first two equations you typed are the ones we are using for solving for x and y? The other one off to the side is maybe what you came up with and you're asking if it's right? Anyway, I will go with the assumption that we are working with the first two and we need to do this with the substitution method. First we can solve the second equation for x. To do this, you subtract 6y from both sides to give 18-6y=x. Now that we have some sort of value for x we can use that value in the other equation where we see the x. This other equation then becomes 2(18-6y) - 3y = -24. See, now we have only one variable instead of 2. Yay! We must distribute the 2 into the (18-6y). That gives us 36-12y. Our equation now looks like this: 36-12y-3y=-24. Now combine like terms to get 36-15y=-24. Isolate the -15y by subtracting the 36 from both sides leaving us with -15y=-60. Now divide both sides by -15 and you get y = 4. Now that you know y = 4, you can sub in 4 where you see a y in either one of your original equations. x + 6(4) = 18 and solve for x. x = what?

OpenStudy (anonymous):

-6?